OpenStudy (anonymous):

How do you factor x^4 + 1 into real factors?

OpenStudy (bahrom7893):

u can't

OpenStudy (bahrom7893):

not over reals.. well depends on what that equals

OpenStudy (lgbasallote):

irreducible

OpenStudy (nikvist):

\[x^4+1=x^4+2x^2+1-2x^2=(x^2+1)^2-(x\sqrt{2})^2=\]\[=(x^2-x\sqrt{2}+1)(x^2+x\sqrt{2}+1)\]